OpenStudy (anonymous):

through (3, -9) parallel to 2x+3y=-12

OpenStudy (anonymous):

Set function in slope-intercept form, then use the inverse of the slope and the point to get the equation of the new line.

OpenStudy (anonymous):

Oops, sorry. You said parallel, so just use the same slope.

OpenStudy (anonymous):

i did that and got y=-2/3+4

OpenStudy (anonymous):

they want it in Ax+by=c

OpenStudy (anonymous):

Just move the term with the x to the same side as the y term.

OpenStudy (anonymous):

i am not suppose to even have a fraction as the slope!

OpenStudy (mathlegend):

@heatzero do you know what the answer is?

OpenStudy (anonymous):

@mathlegend i know i have 4 choices, it's multiple choice

OpenStudy (mathlegend):

Are all the slopes 2x?

OpenStudy (anonymous):

no some are -3y

OpenStudy (mathlegend):

slope would be the one with the x

OpenStudy (anonymous):

its written in standard form

OpenStudy (mathlegend):

Tell me is one of the answers... 2x+3y=-7

OpenStudy (anonymous):

no

OpenStudy (mathlegend):

ok

OpenStudy (anonymous):

I have 2x+3y=-9

OpenStudy (anonymous):

Is that even an option?

OpenStudy (anonymous):

@rudy, no

OpenStudy (anonymous):

@MathLegend no

OpenStudy (anonymous):

You said: "i am not suppose to even have a fraction as the slope!" Why do you say that? Also, if you write (3,-9) into the equation I gave you, it proves valid, also, it has the same slope as the original, so I'm having trouble accepting I'm wrong. Not that it's not possible, I just can't find an error.

OpenStudy (anonymous):

@rudy a fraction isn't one of the choices

OpenStudy (anonymous):

@ rudy, oops my mistake a -9 is one of the choices, however i multiplied everything by 3 to cancel out the fraction and got 21

OpenStudy (anonymous):

From an equation in standard form, the formula for the slope is -A/B Coefficient of x, divided by y coefficient and with negative sign. It is a fraction.

OpenStudy (anonymous):

@ rudy to be clearer the choice is: -3x+2y=-9 and that's not parallel

OpenStudy (anonymous):

Really sorry. Wife's already in the car. Gotta go. Sorry I can't help further. Good luck!

OpenStudy (anonymous):

ok

OpenStudy (mathlegend):

What answers have 2x in them?

OpenStudy (anonymous):

2x+3y =33, 2x+3y=-21

OpenStudy (mathlegend):

@Hero I apologize for requesting you. However, Rudy and I have tried to help and both of us could not figure this out. I was hoping maybe you could try to help here?

hero (hero):

Firstly, it is easier to work with a problem like this if it is in the form y = mx + b, so put it in that form. Secondly, two lines are parallel if they have the same slope, so the slope of the given line is the same as the slope of the line you're trying to find. Thirdly, use y = mx + b, to insert the slope and the given point into the line, then solve for b. Lastly, rewrite the line you need to find in form y = mx + b, then manipulate it back to standard form ax + by = c.

OpenStudy (mathlegend):

I had got y= 2/3x - 4

hero (hero):

I'm pretty sure the slope is negative

OpenStudy (mathlegend):

y=-2/3x-4

hero (hero):

You don't know what the value of b is

hero (hero):

for the new line yet.

hero (hero):

The b values will not be the same for both lines.

OpenStudy (mathlegend):

y=-2/3x-4 So I'll change this into the standard form

OpenStudy (anonymous):

2x+3y=-21, b was -4

OpenStudy (mathlegend):

3y=2x-4

OpenStudy (mathlegend):

-2x+3y=-4

hero (hero):

I'm going to verify this b = -4 that you claim

OpenStudy (anonymous):

ok

OpenStudy (mathlegend):

I did something wrong.... My slope is wrong.

hero (hero):

Your slope is not wrong, but your b value is wrong

hero (hero):

b is not -4 for the NEW line

hero (hero):

I keep telling both of you the b value for the new line will not be the same.

OpenStudy (mathlegend):

I don't know how to find b for this new line. I changed what they gave to slope intercept form.

hero (hero):

Let me show you the steps

OpenStudy (anonymous):

3 or 2

hero (hero):

Firstly, it is easier to work with a problem like this if it is in the form y = mx + b, so put it in that form: \(y = -\frac{2}{3}x -4\) Secondly, two lines are parallel if they have the same slope, so the slope of the given line is the same as the slope of the line you're trying to find. We don't know the value of b yet for this line: \(y = -\frac{2}{3}x + b\) Thirdly, use y = mx + b, to `insert the slope and the given point into the line, then solve for b`: \(-9 = -\frac{2}{3}(3) + b \\-9 = -2 + b \\-9 + 2 = b \\-7 = b\) Lastly, rewrite the line you need to find in form y = mx + b, then manipulate it back to standard form ax + by = c. \(y = -\frac{2}{3}x - 7\) \(\frac{2}{3}x + y = -7\)

hero (hero):

Any questions?

OpenStudy (anonymous):

it's not an option, -7

OpenStudy (mathlegend):

y=-2/3x-7 multiply both sides by 3

OpenStudy (mathlegend):

3y=-2x-21 add 2x to both sides 2x+3y= -21

OpenStudy (mathlegend):

Thank You so much @Hero I never encountered a problem such as this. You are a genius!

OpenStudy (anonymous):

@mathllegend thats the one i picked

OpenStudy (mathlegend):

Yes, but you did not know how to get the answer..

hero (hero):

Good call on multiplying both sides by 3

OpenStudy (mathlegend):

:)

OpenStudy (anonymous):

well i'm gonna close it now, thanks everyone

hero (hero):

@MathLegend, I'm not a genius. This is all basic algebra any math major should know.